In a poll, respondents were asked whether they had ever been in a car accident. 395 respondents indicated that they had been in a car accident and 143 respondents said that they had not been in a car accident. If one of these respondents is randomly selected, what is the probability of getting someone who has been in a car accident?
Answer by stanbon(75887)

In a poll, respondents were asked whether they had ever been in a car accident. 395 respondents indicated that they had been in a car accident and 143 respondents said that they had not been in a car accident. If one of these respondents is randomly selected, what is the probability of getting someone who has been in a car accident?
--------------
P(accident) = 143/(143+395) = 143/538
--------------
Cheers,
Stan H.
-----------